### v1

Abstract
We propose a novel framework that integrates medical image segmentation and point cloud upsampling for accurate shape reconstruction of pelvic models. Using the SAM-Med3D model for segmentation and a point cloud upsampling network trained on the MedShapeNet dataset, our method transforms sparse medical imaging data into high-resolution 3D bone models. This framework leverages prior knowledge of anatomical shapes, achieving smoother and more complete reconstructions. Quantitative evaluations using metrics such as Chamfer Distance etc, demonstrate the effectiveness of the point cloud upsampling in pelvic model. Our approach offers potential applications in reconstructing other skeletal structures, providing a robust solution for medical image analysis and statistical shape modeling.
Keywords:
Statistical Shape Modeling Point Cloud Upsampling Bone Shape Pelvic Segmentation.
1 Introduction
Figure 1: The flowchart of Point Cloud Upsampling as Statistical Shape Model proposed by us requires training a point cloud upsampling network to learn the distribution of bone shapes before inputting volumetric medical images. Subsequently, the image is input into the segmentation model, converted into a point cloud, and upsampled to reconstruct the complete model.
Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions. It is undeniable that U-Net[22] and its various deep learning segmentation models based on the U-Net architecture[7, 23] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[15, 34]. Additionally, various Segment Anything Model (SAM) approaches for medical imaging[12, 26, 18] are becoming increasingly prevalent. Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], automatic recognition[4], and medical diagnosis[11]. Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[24], or for in-vivo bone motion tracking[3].
Creating ground truth for SSM differs from the masks used in semantic segmentation. Training an SSM requires manually identifying shape landmarks to infer statistical data. To build these shapes, a template grid is gradually deformed to segment training images and match the target boundaries, such as those of organs or bones.
To evaluate new patient scans using SSM, the new scan must be segmented and aligned, and the SSM must be optimized for the entire shape population. In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs. Moreover, deep learning methods, inspired by renowned shape benchmarks in computer vision (CV) such as ShapeNet[6], have further developed MedShapeNet[14], which is based on real patient imaging data and instruments.
Building on this, our work focuses on the Pelvic1k dataset[16] with the goal of learning realistic bone shape features through point cloud upsampling. We sample the Pelvic model from MedShapeNet as a point cloud, allowing the upsampling model to learn these features. We then construct a framework that employs the SAM-Med3D model to segment medical images, converts the segmented voxel results into point clouds, and applies the upsampling model, which has prior knowledge of real data, to perform point cloud upsampling. This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy. The integration of medical image segmentation and point cloud upsampling within this framework enables SSM to describe the average shape and facilitates segmentation and model reconstruction.
In summary, the key contributions of our work are as follows:
1) We demonstrate the necessity of learning prior knowledge from MedShapeNet by sampling the pelvic model as point clouds and constructing a baseline using two SOTA point cloud upsampling models, comparing the results with those based on PU1k.
2) We apply the SAM-Med3D model to the Pelvic1k dataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model. This process results in smoother, more complete 3D shape reconstructions.
3) The proposed framework, which integrates medical image segmentation and point cloud upsampling, has the potential to be transferred to various other skeletal models and organs.
2 Related Works
2.1 Point Cloud Processing
Since the introduction of PointNet[19, 20], this direct processing avoids the loss of context and complex steps involved in the data conversion process. Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details. With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks. For example, in point cloud processing tasks, PCN[31] first employs a similar encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input. Later, ASFM-Net[27] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
Besides, PU-Net[30] is a pioneering work that introduces CNNs for point cloud upsampling. Subsequently, Grad-PU[10] similarly provides flexible upsampling at different scales. PUCRN[8] optimized point cloud generation and performance through a multi-stage structure and Transformers[25, 33], especially in sparse or missing regions.
In medical image applications, point cloud completion was first explored by Adams and Elhabian[1], who used [31] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds. And Gu et al. [9] proposed a new approach for 3D reconstruction of individual X-ray images, using point cloud completion to reconstruct real bones from unseen CT images. Later, Yassin et al.[28] proposed a multimodal foundation model for medical point cloud completion based on MedShapeNet[14]. This model is designed to handle incomplete 3D point cloud data and reconstruct the complete shape of various medical structures. By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction. Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
2.2 Segment Anything
The Segment Anything Model (SAM)[12] was trained on a dataset of 11 million images and over 1 billion masks. This design allows the model to be enabled zero-shot transfer to new image distributions and tasks. Due to its speed and versatility, SAM’s zero-shot capabilities have led to its use in medical image segmentation tasks, giving rise to MedSAM[18]. The MedSAM model[18] was developed on a large-scale medical image dataset containing 1.5 million images and masks, covering various imaging modalities and tumor types. It offers better accuracy and robustness than modality-specific models. However, MedSAM[18] only trained by organs and skin, including endoscopic images, and lacks 3D medical images. Later, SAM-Med3D[26] was introduced, focusing on volumetric medical datasets for training and producing satisfactory 3D results. Notably, its training targets include not only organs, skin, and cells but also bones. It is worth mentioning that SSM represents shape variation through statistical analysis of anatomical variations in shape populations generated from medical images, and it has applications in multiple medical fields, including facial recognition and skull analysis[5]. This drives our objective of combining deep learning methods for medical image segmentation and point cloud upsampling, enabling SSM modeling and shape variation.
3 Methodology
3.1 Proposed Framework
In this paper, we propose a framework called Point Cloud Upsampling as Statistical Shape Model (PUSSM), which combines deep learning-based medical image segmentation with point cloud upsampling. Specifically, we first use a real medical shape dataset based on MedShapeNet and apply it to a point cloud upsampling network. This enables us to learn prior knowledge of real-world medical shapes, allowing the upsampling of point clouds from sparse to dense points in anatomical shapes, as shown in Figure 1 (a).
Next, as illustrated in Figure 1 (b), volumetric medical images are input into the pre-trained SAM-Med3D model to obtain segmentation voxel results, which are then converted into point clouds using spatial coordinates. These sparse point clouds are then fed into the point cloud upsampling network to generate a more refined point cloud. Finally, surface reconstruction is performed on the refined point cloud to obtain a denser, higher-resolution volumetric medical model.
The entire process follows the concept of Statistical Shape Models (SSM)[13], where accurate anatomical shapes are learned from the dataset (prior knowledge), and segmentation is used for modeling. Ultimately, similar to visual methods, we achieve segmentation and modeling through SSM’s description of the average shape. However, our network is not end-to-end; instead, it consists of two separate networks: one for medical image segmentation and the other for point cloud upsampling.
3.2 Evaluation Metrics
To evaluate our point cloud upsampling network, we used three key metrics: Chamfer Distance (CD), Hausdorff Distance (HD), and Point-to-Surface Distance (P2F).
Chamfer Distance (CD) measures the average of the nearest-neighbor distances between two point sets.
Hausdorff Distance (HD) captures the maximum distance between points in one set and their closest point in the other set, thus highlighting the worst-case mismatch.
Point-to-Surface Distance (P2F) evaluates the shortest distance from upsampled points to the surface of the reference model.
4 Experiments
4.1 Settings
4.1.1 Datasets
In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[14] for point cloud upsampling to learn shape prior knowledge. We then visualized subject’s result based on Pelvic1k[16], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
For training and testing the upsampling network, we used the MedShape Pelvic dataset based on MedShapeNet[14], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[29]. The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples. Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
4.1.2 Comparison Methods
To verify the effectiveness of our proposed theory in point cloud upsampling as statistical shape model, comparing it with two advanced algorithms: Grad-Pu[10] that realized any ratio upsampling and PUCRN[8] that is the SOTA model.
For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment. And we test the result based on training in MedShape Pelvic[14] and PU1k dataset[21], which is a synthetic dataset from 3D models. This comparison with identical parameter settings would be presented in ablation: Section 4.3.
4.1.3 Implementation Details
Experiments were developed using the PyTorch framework on the Ubuntu 22.04 system. We used an NVIDIA L20 GPU with 48GB of graphics memory and an Intel(R) Xeon(R) Platinum 8457C host with 100GB of RAM. The network was trained over 100 epochs with a batch size of 64. The initial learning rate was set to 0.001, with a decay rate of 0.05.
4.2 Upsampling Results
In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShape Pelvic dataset in 4×4$\times$4 × upsampling. The quantitative results for the dataset in Table 1 show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
This result demonstrates the effectiveness of PUCRN in CD, HD, and P2F, showcasing its superior capability in learning the shape prior knowledge in pelvic. The 3 stage network refinement structure allow it could minimizing nearest-neighbor distance differences and reducing maximum point-to-point distance errors.
Additionally, noticed that Grad-PU with very high HD. This may be caused by the input pelvic training data. This reflects that Grad PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 2 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2. Grad PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set. However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect. The results show that PUCRN still has good performance.
Table 1: Performance comparison of two methods with MedShape Pelvic across input scales. Among them, the unit is $10^{-3}$.
Methods 512 input 1024 input 2048 input 4096 input
[…]
Grad-PU 5.643 124.639 12.438 3.682 82.969 9.779 2.477 49.451 4.852 1.690 28.871 3.172
PUCRN 3.104 7.806 6.150 3.946 16.635 6.436 1.839 5.709 4.066 1.274 5.087 2.719
4.3 Ablation Study
To verify the effectiveness and necessity of training in MedShape Pelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShape Pelvic[14] and PU1k dataset[21] have been tested. Among them, MedShape Pelvic has 50496 training samples and PU1k has 69000 training samples. This avoids the difference in learning effectiveness caused by the large difference in the number of training samples. The result shows in Table 3
Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024. Among them, the unit is $10^{-3}$.
Methods 512 to 8192 1024 to 16384
[…]
Grad-PU 5.827 124.640 13.117 3.422 84.825 10.329
PUCRN 3.586 9.984 6.824 2.870 18.899 5.872
As a result, training on the MedShape Pelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad PU’s model on the PU1k dataset performed better than MedShape Pelvic’s test set. These reasons are multifaceted, and Grad PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShape Pelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets.
Methods 512 input 1024 input 2048 input 4096 input
[…]
Grad-PU(MedShapeNet) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
PUCRN(MedShapeNet) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
Grad-PU(PU1k) 5.411 15.551 3.353 12.280 2.256 8.677 1.513 7.872
PUCRN(PU1k) 3.749 10.065 2.902 9.955 1.995 6.134 1.379 5.253
4.4 Visualization
In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM. In this visualization section, due to the lack of ground truth, we can only choose to display the reconstruction effect. If we want to make quantitative comparisons, it involves the need for No Reference Mesh Quality Assessment. However, mostly 3D quality assessment focusing on point cloud instead of mesh [32].
Figure 2: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
Now we input the Pelvic1k dataset into SAM Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[17]. As shows in Figure 2, compared with the surface obtained by directly segmenting voxels through marching cubes (Pseudo Ground Truth), our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN.
5 Conclusion
This study introduced a framework integrating point cloud upsampling with statistical shape modeling to improve high-resolution pelvic model reconstruction. By combining SAM-Med3D segmentation with upsampling networks, we transformed sparse voxel data into dense point clouds and reconstructed detailed surfaces using the Marching Cube algorithm. Experiments demonstrated the necessity of upsampling and confirmed the framework’s effectiveness in achieving smoother and more complete reconstructions.
Future work will extend this framework to other anatomical structures and incorporate no reference quality assessment for objective evaluations.
{credits}
5.0.1 Acknowledgements
This research was supported by the National Natural Science Foundation of China under Grant 61773164.
[…]
The Department of Rehabilitation Science, The Hong Kong Polytechnic University, has not financial support, and there are no potential conflicts of interest to disclose.

### v3

@@
-
+PUSSM: Point Cloud Upsampling as Implicit Statistical Shape Model
 Abstract
-We propose a novel framework that integrates medical image segmentation and point cloud upsampling for accurate shape reconstruction of pelvic models.
-Using the SAM-Med3D model for segmentation and a point cloud upsampling network trained on the MedShapeNet dataset, our method transforms sparse medical imaging data into high-resolution 3D bone models.
-This framework leverages prior knowledge of anatomical shapes, achieving smoother and more complete reconstructions.
-Quantitative evaluations using metrics such as Chamfer Distance etc, demonstrate the effectiveness of the point cloud upsampling in pelvic model.
-Our approach offers potential applications in reconstructing other skeletal structures, providing a robust solution for medical image analysis and statistical shape modeling.
+This paper proposes a framework for high-fidelity reconstruction of pelvic structures by integrating medical image segmentation and point cloud upsampling.
+By point cloud upsampling to learn shape priors from MedShapePelvic without requiring landmarks or PCA, our method functions as an implicit statistical shape model.
+Evaluations on Pelvic1k show significant improvements in surface quality and anatomical accuracy.
+This approach is generalizable and applicable to other skeletal regions.
 Keywords:
@@
 1 Introduction
-Figure 1: The flowchart of Point Cloud Upsampling as Statistical Shape Model proposed by us requires training a point cloud upsampling network to learn the distribution of bone shapes before inputting volumetric medical images.
-Subsequently, the image is input into the segmentation model, converted into a point cloud, and upsampled to reconstruct the complete model.
 Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions.
-It is undeniable that U-Net[22] and its various deep learning segmentation models based on the U-Net architecture[7, 23] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[15, 34].
+It is undeniable that U-Net[25] and its various deep learning segmentation models based on the U-Net architecture[6, 26] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[36].
 Additionally, various Segment Anything Model (SAM) approaches for medical imaging[15, 28, 21] are becoming increasingly prevalent.
-Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], automatic recognition[4], and medical diagnosis[11].
+Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], and medical diagnosis[14].
 Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[27], or for in-vivo bone motion tracking[3].
+Figure 1: The proposed Point Cloud Upsampling as Statistical Shape Model involves training a point cloud upsampling network to learn bone shape distributions from volumetric medical images.
+The image is then segmented into a point cloud, upsampled, and reconstructed into a complete model.
 Creating ground truth for SSM differs from the masks used in semantic segmentation.
@@
 In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs.
-Moreover, deep learning methods, inspired by renowned shape benchmarks in computer vision (CV) such as ShapeNet[6], have further developed MedShapeNet[14], which is based on real patient imaging data and instruments.
-Building on this, our work focuses on the Pelvic1k dataset[16] with the goal of learning realistic bone shape features through point cloud upsampling.
-We sample the Pelvic model from MedShapeNet as a point cloud, allowing the upsampling model to learn these features.
-We then construct a framework that employs the SAM-Med3D model to segment medical images, converts the segmented voxel results into point clouds, and applies the upsampling model, which has prior knowledge of real data, to perform point cloud upsampling.
+Moreover, deep learning methods, inspired by renowned shape benchmarks in Computer Vision (CV) have further developed MedShapeNet[16], which is based on real patient imaging data and instruments.
+However, classical SSM construction typically involves manual annotation of anatomical landmarks and tedious correspondence alignment, which may limit scalability and automation in large datasets.
+In contrast, our approach explores a data-driven alternative to SSM, where prior anatomical knowledge is learned implicitly through point cloud upsampling networks trained on large-scale 3D medical datasets, which is on the Pelvic model of MedShapeNet dataset (MedShapePelvic)[16] with the goal of learning realistic bone shape features through point cloud upsampling.
+Instead of modeling shape variation explicitly via PCA, our method leverages learned feature distributions from dense anatomical shapes to enhance the reconstruction of sparse pelvic scans.
+This shift from explicit modeling to learned priors enables more flexible and automatic shape analysis, especially when integrated with modern segmentation tools such as SAM-Med3D.
 This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy.
@@
 In summary, the key contributions of our work are as follows:
-1) We demonstrate the necessity of learning prior knowledge from MedShapeNet by sampling the pelvic model as point clouds and constructing a baseline using two SOTA point cloud upsampling models, comparing the results with those based on PU1k.
-2) We apply the SAM-Med3D model to the Pelvic1k dataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model.
+1) We demonstrate the necessity of learning prior knowledge from MedShapePelvic by sampling the pelvic model as point clouds and constructing a baseline using four SOTA point cloud upsampling models, comparing the results with those based on PU1k.
+2) We apply the SAM-Med3D model to the Pelvic1k subdataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model.
 This process results in smoother, more complete 3D shape reconstructions.
@@
 2.1 Point Cloud Processing
+2.1.1 In 3D Object
 Since the introduction of PointNet[22, 23], this direct processing avoids the loss of context and complex steps involved in the data conversion process.
-Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details.
+Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details[18, 17].
 With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks.
@@
 Besides, PU-Net[32] is a pioneering work that introduces CNNs for point cloud upsampling.
-Subsequently, Grad-PU[10] similarly provides flexible upsampling at different scales.
-PUCRN[8] optimized point cloud generation and performance through a multi-stage structure and Transformers[25, 33], especially in sparse or missing regions.
-In medical image applications, point cloud completion was first explored by Adams and Elhabian[1], who used [31] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds.
+Subsequently, PU-GCN[24] employed a graph-based network construction, achieving good upsampling performance.
+Grad-PU[13] similarly provides flexible upsampling at different scales.
+PUCRN[9] optimized point cloud generation and performance through a multi-stage structure and Transformers[35], especially in sparse or missing regions.
+However, these methods are often trained on synthetic datasets (e.g., PU1k dataset[24] proposed by PU-GCN) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling.
+Moreover, they typically focus on object-level CAD models and are not directly validated on medical segmentation pipelines.
+2.1.2 In Medical Image
+For medical image applications, Chen et al.[7] proposed a learning grid to point cloud deformable shape registration framework from multiple 2D image slices, which can achieve accurate 3D grid reconstruction in real-time in the absence of data and sparse annotations.
+Point cloud completion was first explored by Adams and Elhabian[1], who used [33] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds.
 And Gu et al.
@@
 By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction.
-Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
+Zhang et al.[34] proposed a hierarchical feature learning framework based on the State Space Model for medical point cloud understanding, which will aid in anatomical classification, completion, and segmentation.Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
+In contrast to these works, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling.
+Furthermore, by training on realistic datasets, MedShapePelvic, and test on Pelvic1k subdataset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
 2.2 Segment Anything
@@
 It offers better accuracy and robustness than modality-specific models.
-However, MedSAM[18] only trained by organs and skin, including endoscopic images, and lacks 3D medical images.
+However, MedSAM[21] only trained by organs and skin, including endoscopic images, and lacks 3D medical images, because they focus on interactive in 2D images.
 Later, SAM-Med3D[28] was introduced, focusing on volumetric medical datasets for training and producing satisfactory 3D results.
@@
 This drives our objective of combining deep learning methods for medical image segmentation and point cloud upsampling, enabling SSM modeling and shape variation.
+Ergo, our study leverages SAM-Med3D as a frontend for downstream statistical shape modeling.
+By combining its segmentation output with point cloud upsampling, we demonstrate that deep segmentation models can serve as a viable input source for high-fidelity shape reconstruction.
 3 Methodology
-3.1 Proposed Framework
-In this paper, we propose a framework called Point Cloud Upsampling as Statistical Shape Model (PUSSM), which combines deep learning-based medical image segmentation with point cloud upsampling.
-Specifically, we first use a real medical shape dataset based on MedShapeNet and apply it to a point cloud upsampling network.
-This enables us to learn prior knowledge of real-world medical shapes, allowing the upsampling of point clouds from sparse to dense points in anatomical shapes, as shown in Figure 1 (a).
-Next, as illustrated in Figure 1 (b), volumetric medical images are input into the pre-trained SAM-Med3D model to obtain segmentation voxel results, which are then converted into point clouds using spatial coordinates.
-These sparse point clouds are then fed into the point cloud upsampling network to generate a more refined point cloud.
-Finally, surface reconstruction is performed on the refined point cloud to obtain a denser, higher-resolution volumetric medical model.
-The entire process follows the concept of Statistical Shape Models (SSM)[13], where accurate anatomical shapes are learned from the dataset (prior knowledge), and segmentation is used for modeling.
-Ultimately, similar to visual methods, we achieve segmentation and modeling through SSM’s description of the average shape.
-However, our network is not end-to-end; instead, it consists of two separate networks: one for medical image segmentation and the other for point cloud upsampling.
+3.1 Overview of the Proposed Framework
+We propose a novel framework called Point Cloud Upsampling as Implicit Statistical Shape Model (PUSSM), which integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction.
+As shown in Figure 1(a), the core idea is to learn an implicit shape prior from real anatomical data through a deep upsampling network, allowing sparse point clouds to be restored to dense representations in an anatomically plausible way.
+The pipeline consists of two sequential modules.
+First, we apply the SAM-Med3D model to volumetric CT data to obtain voxel-wise segmentation masks.
+These masks are converted into sparse point clouds using spatial coordinates.
+Then, as illustrated in Figure 1(b), the sparse point cloud is passed into a pretrained point cloud upsampling network, which has been trained on MedShapePelvic samples.
+Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes.
+Figure 2: Distributional interpretation of implicit statistical shape modeling.
+The segmentation-derived point cloud (left) deviates from the real anatomical distribution.
+Our network $f_{\theta}$ learns to map the sparse inputs toward the ground-truth distribution through point cloud upsampling.
+3.2 Implicit Statistical Shape Modeling
+Unlike classical Statistical Shape Models (SSMs), which explicitly encode anatomical variation via PCA over aligned meshes, our framework learns an implicit statistical shape model through data-driven training.
+Specifically, let $X\subset\mathbb{R}^{3}$ denote a sparse point cloud from SAM-Med3D, and let $Y\subset\mathbb{R}^{3}$ be the corresponding ground truth dense point cloud from MedShapePelvic.
+We train a deep function $f_{\theta}:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{M\times 3}$ to minimize the geometric discrepancy between $f_{\theta}(X)$ and $Y$, using a distance metric $D(\cdot,\cdot)$ such as the Chamfer Distance:
+$\min_{\theta}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[D(f_{\theta}(X),Y)\right]$
+This process embeds the shape prior directly into the network weights $\theta$, enabling the model to generalize to new, noisy, or incomplete segmentation results.
+Thus, $f_{\theta}$ serves as an implicit statistical shape model learned directly from anatomical data—requiring no landmark alignment, shape vectors, or PCA decomposition.
+Figure 2 further illustrates this concept from a distributional perspective, showing how our model maps the segmentation-derived distribution toward the ground-truth anatomical distribution via upsampling.
+Figure 3: Conceptual comparison between classical SSM and our implicit PUSSM framework across key dimensions.
+3.3 Comparison with Classical SSMs
+To highlight the differences between our method and classical SSMs, Figure 3 provides a conceptual comparison across three key dimensions: shape representation, prior data source, and learning strategy.
+Traditional SSMs rely on manually constructed shape vectors and PCA decomposition of landmark-aligned meshes.
+In contrast, our method encodes shape priors directly through end-to-end learning from dense mesh data, enabling greater flexibility and broader applicability.
 3.4 Evaluation Metrics
-To evaluate our point cloud upsampling network, we used three key metrics: Chamfer Distance (CD), Hausdorff Distance (HD), and Point-to-Surface Distance (P2F).
+3.4.1 For Point Cloud Upsampling
+To evaluate the point cloud upsampling network, we used three key metrics:
 Chamfer Distance (CD) measures the average of the nearest-neighbor distances between two point sets.
+Let $P$ and $Q$ be the two sets of points (the upsampled point cloud and the ground truth point cloud, respectively), the Chamfer Distance is defined as:
+$\text{CD}(P,Q)=\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|p-q\|^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|q-p\|^{2}$
 Hausdorff Distance (HD) captures the maximum distance between points in one set and their closest point in the other set, thus highlighting the worst-case mismatch.
+The Hausdorff distance between sets $P$ and $Q$ is defined as:
+$\text{HD}(P,Q)=\max\left(\max_{p\in P}\min_{q\in Q}\|p-q\|,\max_{q\in Q}\min_{p\in P}\|q-p\|\right)$
 Point-to-Surface Distance (P2F) evaluates the shortest distance from upsampled points to the surface of the reference model.
+Given a set of upsampled points $P$ and a reference surface $S$, the Point-to-Surface Distance is defined as:
+[…]
+3.4.2 For Surface Reconstruction
+To evaluate the surface reconstruction of our framework, we used three key metrics: Area-Length Ratio, Manifoldness Rate and Connected Component Discrepancy[11, 4].
+Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh.
+This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality.
+It is defined as:
+Manifoldness Rate (MR) quantifies the proportion of edges in a mesh that satisfy manifold conditions.
+An edge is considered manifold if it is shared by exactly two faces.
+A rate approaching 1 indicates a well-structured mesh suitable for operations like Boolean computations and 3D printing.
+Conversely, a lower rate highlights the presence of non-manifold edges, which may necessitate mesh repair or refinement.
+It is defined as:
+Manifoldness Rate=Number of manifold edgesTotal number of $\text{Manifoldness Rate}=\frac{\text{Number of manifold edges}}{\text{Total number of edges}}$
+Connected Component Discrepancy (CC Diff.) assesses the topological connectedness of a mesh by comparing the observed number of connected components to the expected number.
+A discrepancy of 0 signifies that the mesh has the anticipated number of connected components, implying structural integrity.
+It is defined as:
+[…]
+Those three indicators to evaluate the geometric quality of reconstructed meshes.
+Besides, to evaluate the fidelity of the reconstructed meshes against the ground truth, we employ several widely recognized metrics[8]: Chamfer Distance (CD), F-Score (F1), Normal Consistency (NC), Edge Chamfer Distance (ECD), and Edge F-Score (EF1).
 4 Experiments
@@
 4.1.1 Datasets
-In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[14] for point cloud upsampling to learn shape prior knowledge.
-We then visualized subject’s result based on Pelvic1k[16], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
+In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[16] as MedShapePlevic for point cloud upsampling to learn shape prior knowledge.
+We then visualized Quantitative analyzed result based on Pelvic1k[19], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
+Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g.
+subdatasets 3, 4, 6).
 For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[16], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[31].
@@
 4.1.2 Comparison Methods
-To verify the effectiveness of our proposed theory in point cloud upsampling as statistical shape model, comparing it with two advanced algorithms: Grad-Pu[10] that realized any ratio upsampling and PUCRN[8] that is the SOTA model.
+In order to verify the effectiveness of our proposed theory, point cloud upsampling is the first part.
+We built a baseline based on MPU[31], PU-GCN[24], Grad-Pu[13] and PUCRN[9].
+In more experiments, only by comparing it with two advanced algorithms: Grad-Pu that realized any ratio upsampling and PUCRN that is the SOTA model.
 For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment.
@@
 This may be caused by the input pelvic training data.
-This reflects that Grad PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 2 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
+This reflects that Grad-PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 4 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
 Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2.
-Grad PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set.
+Grad-PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set.
 However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect.
 The results show that PUCRN still has good performance.
-Table 1: Performance comparison of two methods with MedShape Pelvic across input scales. Among them, the unit is $10^{-3}$.
-Methods 512 input 1024 input 2048 input 4096 input
+Table 1: Performance comparison of two methods with MedShapePelvic across input scales.
+Method 512 input 1024 input 2048 input 4096 input
 […]
+MPU 13.050 11.971 10.604 8.507 26.088 8.759 5.718 9.421 8.558 3.906 8.768 7.556
+PU-GCN 4.191 9.942 8.553 5.326 21.284 7.329 2.483 7.296 4.652 1.719 6.519 3.671
 Grad-PU 5.643 124.639 12.438 3.682 82.969 9.779 2.477 49.451 4.852 1.690 28.871 3.172
 PUCRN 3.104 7.806 6.150 3.946 16.635 6.436 1.839 5.709 4.066 1.274 5.087 2.719
-4.3 Ablation Study
-To verify the effectiveness and necessity of training in MedShape Pelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShape Pelvic[14] and PU1k dataset[21] have been tested.
+4.3 Ablation Studies
+To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShapePelvic and PU1k dataset[24] have been tested.
 Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples.
@@
 The result shows in Table 3
-Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024. Among them, the unit is $10^{-3}$.
-Methods 512 to 8192 1024 to 16384
+Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024.
+Method 512 to 8192 1024 to 16384
 […]
@@
 This is very important.
-However, Grad PU’s model on the PU1k dataset performed better than MedShape Pelvic’s test set.
-These reasons are multifaceted, and Grad PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShape Pelvic obtained from real volume medical images.
-Table 3: Pelvic point cloud upsampling results in two different train sets.
-Methods 512 input 1024 input 2048 input 4096 input
+However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set.
+These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
+Table 3: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
+Method 512 input 1024 input 2048 input 4096 input
 […]
-Grad-PU(MedShapeNet) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
-PUCRN(MedShapeNet) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
+Grad-PU(MedShapePelvic) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
+PUCRN(MedShapePelvic) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
 Grad-PU(PU1k) 5.411 15.551 3.353 12.280 2.256 8.677 1.513 7.872
 PUCRN(PU1k) 3.749 10.065 2.902 9.955 1.995 6.134 1.379 5.253
-4.4 Visualization
+Figure 4: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
+4.4 Reconstruction Visualizations
 In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM.
-In this visualization section, due to the lack of ground truth, we can only choose to display the reconstruction effect.
-If we want to make quantitative comparisons, it involves the need for No Reference Mesh Quality Assessment.
-However, mostly 3D quality assessment focusing on point cloud instead of mesh [32].
-Figure 2: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
-Now we input the Pelvic1k dataset into SAM Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds.
+The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds.
 The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling.
-The model surface is reconstructed using the Voxel Grid with Marching Cubes[17].
-As shows in Figure 2, compared with the surface obtained by directly segmenting voxels through marching cubes (Pseudo Ground Truth), our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN.
+The model surface is reconstructed using the Voxel Grid with Marching Cubes[20], where voxel size is $1.5$.
+As shows in Figure 4, compared with the surface of ground truth, and obtained by directly segmenting voxels through marching cubes.
+Our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN.
+And there are no obvious contours of layer after reconstructed.
+4.5 Reconstruction Evaluations
+To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in Section3.4.2 For Surface Reconstruction.
+Table 4: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
+Subdataset 3 Subdataset 4 Subdataset 6
+[…]
+Ground Truth 0.692 100.0 0.26 0.721 100.0 1.68 0.742 100.0 0.66
+Segmented Result (by SAM-Med3D) 0.676 100.0 1.94 0.699 100.0 4.89 0.735 100.0 1.18
+PUSSM (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
+PUSSM (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
+As shows in Table 4.
+Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy.
+Specifically, our method has uniform grid partitioning (as shown in ALR) and eliminates floating artifacts (as shown in CC Diff.).
+This improvement is attributed to the finer points and distribution brought about by point cloud upsampling.
+Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation.
+We used metrics (CD, F1, NC, ECD, EF1) in Section3.4.2 For Surface Reconstruction.
+To ensure a fair and uniform comparison, we employ Poisson disk sampling to generate 16,384 points from each mesh, including the ground truth, segmented results, and upsampled reconstructions.
+This sampling strategy maintains uniform point distribution and adequately captures both global structures and fine details, facilitating reliable metric computation.
+Table 5: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
+Subdataset 3 Subdataset 4 Subdataset 6
+[…]
+Segmented Result (by SAM-Med3D) 8.379 0.214 0.691 7.420 0.408 56.272 0.126 0.711 54.559 0.250 13.997 0.214 0.716 10.724 0.357
+PUSSM (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
+PUSSM (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
+As shows in Table 5.
+Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations.
+For example, on Subdataset 3, PUCRN achieves a significantly lower CD compared to Grad-PU, along with superior edge preservation as reflected by a lower ECD and a higher EF1-score.
+Similar trends are observed on Subdataset 4 and Subdataset 6, where PUCRN achieves the best overall performance in terms of both global surface quality and local edge fidelity.
+These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions.
 5 Conclusion
-This study introduced a framework integrating point cloud upsampling with statistical shape modeling to improve high-resolution pelvic model reconstruction.
-By combining SAM-Med3D segmentation with upsampling networks, we transformed sparse voxel data into dense point clouds and reconstructed detailed surfaces using the Marching Cube algorithm.
-Experiments demonstrated the necessity of upsampling and confirmed the framework’s effectiveness in achieving smoother and more complete reconstructions.
-Future work will extend this framework to other anatomical structures and incorporate no reference quality assessment for objective evaluations.
+In this work, we proposed PUSSM, an implicit statistical shape modeling framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data.
+Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation.
+Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
+Future work will extend this framework to 3D SAM model[15] by points based prompting to integrate it as Chain-of-Thought (CoT)[12] for Segmentation.
 {credits}
-5.0.1 Acknowledgements
-This research was supported by the National Natural Science Foundation of China under Grant 61773164.
-[…]
-The Department of Rehabilitation Science, The Hong Kong Polytechnic University, has not financial support, and there are no potential conflicts of interest to disclose.

### v4

@@
-PUSSM: Point Cloud Upsampling as Implicit Statistical Shape Model
+Med-PU: Point Cloud Upsampling for High-Fidelity 3D Medical Shape Reconstruction
 Abstract
-This paper proposes a framework for high-fidelity reconstruction of pelvic structures by integrating medical image segmentation and point cloud upsampling.
-By point cloud upsampling to learn shape priors from MedShapePelvic without requiring landmarks or PCA, our method functions as an implicit statistical shape model.
-Evaluations on Pelvic1k show significant improvements in surface quality and anatomical accuracy.
-This approach is generalizable and applicable to other skeletal regions.
-Keywords:
-Statistical Shape Modeling Point Cloud Upsampling Bone Shape Pelvic Segmentation.
-1 Introduction
-Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions.
-It is undeniable that U-Net[25] and its various deep learning segmentation models based on the U-Net architecture[6, 26] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[36].
-Additionally, various Segment Anything Model (SAM) approaches for medical imaging[15, 28, 21] are becoming increasingly prevalent.
-Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], and medical diagnosis[14].
-Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[27], or for in-vivo bone motion tracking[3].
-Figure 1: The proposed Point Cloud Upsampling as Statistical Shape Model involves training a point cloud upsampling network to learn bone shape distributions from volumetric medical images.
-The image is then segmented into a point cloud, upsampled, and reconstructed into a complete model.
-Creating ground truth for SSM differs from the masks used in semantic segmentation.
-Training an SSM requires manually identifying shape landmarks to infer statistical data.
-To build these shapes, a template grid is gradually deformed to segment training images and match the target boundaries, such as those of organs or bones.
-To evaluate new patient scans using SSM, the new scan must be segmented and aligned, and the SSM must be optimized for the entire shape population.
-In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs.
-Moreover, deep learning methods, inspired by renowned shape benchmarks in Computer Vision (CV) have further developed MedShapeNet[16], which is based on real patient imaging data and instruments.
-However, classical SSM construction typically involves manual annotation of anatomical landmarks and tedious correspondence alignment, which may limit scalability and automation in large datasets.
-In contrast, our approach explores a data-driven alternative to SSM, where prior anatomical knowledge is learned implicitly through point cloud upsampling networks trained on large-scale 3D medical datasets, which is on the Pelvic model of MedShapeNet dataset (MedShapePelvic)[16] with the goal of learning realistic bone shape features through point cloud upsampling.
-Instead of modeling shape variation explicitly via PCA, our method leverages learned feature distributions from dense anatomical shapes to enhance the reconstruction of sparse pelvic scans.
-This shift from explicit modeling to learned priors enables more flexible and automatic shape analysis, especially when integrated with modern segmentation tools such as SAM-Med3D.
-This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy.
-The integration of medical image segmentation and point cloud upsampling within this framework enables SSM to describe the average shape and facilitates segmentation and model reconstruction.
-In summary, the key contributions of our work are as follows:
-1) We demonstrate the necessity of learning prior knowledge from MedShapePelvic by sampling the pelvic model as point clouds and constructing a baseline using four SOTA point cloud upsampling models, comparing the results with those based on PU1k.
-2) We apply the SAM-Med3D model to the Pelvic1k subdataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model.
-This process results in smoother, more complete 3D shape reconstructions.
-3) The proposed framework, which integrates medical image segmentation and point cloud upsampling, has the potential to be transferred to various other skeletal models and organs.
-2 Related Works
-2.1 Point Cloud Processing
-2.1.1 In 3D Object
-Since the introduction of PointNet[22, 23], this direct processing avoids the loss of context and complex steps involved in the data conversion process.
-Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details[18, 17].
+High-fidelity 3D anatomical reconstruction is a prerequisite for downstream clinical tasks such as preoperative planning, radiotherapy target delineation, and orthopedic implant design.
+We present Med-PU, a knowledge-driven framework that integrates volumetric medical image segmentation with point cloud upsampling for accurate pelvic shape reconstruction.
+Unlike landmark- or PCA-based statistical shape models, Med-PU learns an implicit anatomical prior directly from large-scale 3D shape data, enabling dense completion and refinement from sparse segmentation-derived point sets.
+The pipeline couples SAM-Med3D-based voxel segmentation, point extraction, deep upsampling, and surface reconstruction, yielding smooth and topologically consistent meshes.
+We evaluate Med-PU on pelvic CT datasets (MedShapePelvic for training and Pelvic1k for validation), benchmarking against state-of-the-art upsampling methods using comprehensive geometry and surface metrics.
+Med-PU consistently improves surface quality and anatomical fidelity while reducing artifacts, demonstrating robustness across input densities.
+Although validated on the pelvis, the approach is anatomy-agnostic and applicable to other skeletal regions and organs.
+These results suggest Med-PU as a practical, generalizable tool to bridge segmentation outputs and clinically usable 3D models.
+I Introduction
+High-fidelity 3D anatomical reconstruction is a prerequisite for several downstream clinical tasks, including preoperative planning, radiotherapy target delineation, and orthopedic implant design.
+In routine pipelines, clinicians often begin with voxel-wise segmentation and then derive surface meshes; however, directly converting voxel outputs into smooth, watertight, and anatomically faithful meshes remains challenging due to partial observations, image noise, and inter-subject variability.
+Traditional Statistical Shape Models (SSMs) address variability by aligning training shapes and learning a low-dimensional subspace for population-level morphometrics.
+They have been widely adopted for segmentation, measurement, and diagnosis [1, 2], and can even generate patient-specific bone models from sparse 3D/2D data [3, 4].
+Nevertheless, classical SSM construction depends on manual landmark annotation, explicit correspondence, and PCA, which limits scalability and automation on large medical datasets and can constrain the expressiveness of subtle anatomical details.
+In parallel, modern segmentation networks—U-Net variants [5, 6, 7], nnFormer [8], and SAM-based extensions for medical imaging [9, 10, 11]—deliver robust voxel-level delineations.
+Yet a persistent gap remains between these voxel labels and clinically usable 3D models: meshes reconstructed from discrete masks are often jagged, topologically inconsistent, or over-smoothed, especially under limited resolution or incomplete coverage.
+This motivates learning shape priors directly from data so that sparse or noisy anatomical evidence can be lifted to dense, high-quality surfaces.
+Recent studies are beginning to bridge point cloud learning and medical applications.
+Adams and Elhabian [12] explored unsupervised point-distribution modeling from raw point sets, while MedShapeNet [13] established a large-scale benchmark for medical 3D shape learning.
+These advances indicate that implicit, data-driven priors can complement or replace explicit SSMs in clinical 3D modeling, particularly when combined with strong segmentation front-ends.
+Figure 1: Overview of the proposed pipeline.
+Medical images are first segmented into coarse anatomical masks (e.g., SAM-Med3D), then converted to point clouds, followed by upsampling to reconstruct a complete and high-resolution 3D bone model.
+Unlike classical Statistical Shape Models that require landmarks and PCA, our framework learns shape variability implicitly from dense point cloud data.
+This work
+We propose a data-driven framework for producing clinically amenable meshes from segmentation outputs.
+Instead of explicitly constructing and sampling a statistical shape space, we leverage point cloud upsampling networks trained on large-scale 3D medical shape data to implicitly encode anatomical priors.
+Using the pelvic subset of MedShapeNet (MedShapePelvic) as shape supervision and SAM-Med3D as a volumetric segmentation front-end, our framework learns to complete and refine sparse, segmentation-derived point sets into dense, anatomically faithful surfaces.
+This shift from explicit parametric modeling to learned priors enables flexible, automated reconstruction under varying input densities, helping to bridge the gap between voxel labels and high-quality meshes required in downstream clinical workflows.
+Contributions
+  • A segmentation-to-mesh pipeline with implicit priors.
+We integrate SAM-Med3D-based voxel segmentation with deep point cloud upsampling to learn an implicit anatomical prior that densifies and regularizes segmentation-derived point sets, producing smooth, watertight meshes without manual landmarks or PCA.
+  • Comprehensive evaluation on pelvic CT.
+Trained on MedShapePelvic and validated on Pelvic1k, the proposed method is benchmarked against state-of-the-art upsampling approaches using geometry- and surface-level metrics (Chamfer/Hausdorff distances, point-to-surface error, F-score, normal consistency, and edge-aware measures), demonstrating consistent gains in surface quality and anatomical fidelity.
+  • Generalization potential beyond the pelvis.
+Although evaluated on pelvic bones, the framework is anatomy-agnostic and applicable to other skeletal regions and organs, providing a practical bridge from voxel segmentation to clinically usable 3D models in broader biomedical contexts.
+II Related Works
+II-A Point Cloud Processing
+II-A1 In 3D Object
+Since the introduction of PointNet[14, 15], direct processing of raw point sets avoids the loss of context and complex steps involved in voxelization or multi-view conversion.
+Point cloud completion, in particular, is used to extract global features of 3D models, which can help supplement geometric details[16, 17].
 With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks.
-For example, in point cloud processing tasks, PCN[33] first employs a similar encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input.
+For example, PCN[18] first employs an encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input.
 Later, ASFM-Net[19] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
 Besides, PU-Net[20] is a pioneering work that introduces CNNs for point cloud upsampling.
-Subsequently, PU-GCN[24] employed a graph-based network construction, achieving good upsampling performance.
-Grad-PU[13] similarly provides flexible upsampling at different scales.
-PUCRN[9] optimized point cloud generation and performance through a multi-stage structure and Transformers[35], especially in sparse or missing regions.
-However, these methods are often trained on synthetic datasets (e.g., PU1k dataset[24] proposed by PU-GCN) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling.
+Subsequently, PU-GCN[21] employed a graph-based network, achieving strong upsampling performance.
+Grad-PU[22] provides flexible upsampling at different scales, while PUCRN[23] optimized generation through multi-stage cascades.
+Transformer-based frameworks[24] also improved robustness in sparse or missing regions.
+However, these methods are often trained on synthetic datasets (e.g., PU1k[21]) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling.
 Moreover, they typically focus on object-level CAD models and are not directly validated on medical segmentation pipelines.
-2.1.2 In Medical Image
-For medical image applications, Chen et al.[7] proposed a learning grid to point cloud deformable shape registration framework from multiple 2D image slices, which can achieve accurate 3D grid reconstruction in real-time in the absence of data and sparse annotations.
-Point cloud completion was first explored by Adams and Elhabian[1], who used [33] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds.
-And Gu et al.
-[10] proposed a new approach for 3D reconstruction of individual X-ray images, using point cloud completion to reconstruct real bones from unseen CT images.
-Later, Yassin et al.[30] proposed a multimodal foundation model for medical point cloud completion based on MedShapeNet[16].
-This model is designed to handle incomplete 3D point cloud data and reconstruct the complete shape of various medical structures.
-By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction.
-Zhang et al.[34] proposed a hierarchical feature learning framework based on the State Space Model for medical point cloud understanding, which will aid in anatomical classification, completion, and segmentation.Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
-In contrast to these works, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling.
-Furthermore, by training on realistic datasets, MedShapePelvic, and test on Pelvic1k subdataset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
-2.2 Segment Anything
+II-A2 In Medical Image
+Recently, point cloud learning has been increasingly applied to medical image analysis.
+Chen et al.[25] proposed a learning-based deformable registration framework (MR-Net) to reconstruct 3D cardiac meshes from sparse 2D contours, showing real-time capability even with incomplete annotations.
+Beetz et al.
+developed Point2Mesh-Net, which combines point-based encoders with mesh-based decoders for cardiac surface reconstruction from cine MRI slices[26] , and later extended this line into multi-objective autoencoders for myocardial infarction prediction, enabling explainable outcome modeling from 3D cardiac shapes[27].
+Hu et al.[28] introduced a GAN with tree-structured graph convolutions to generate 3D brain point clouds from a single 2D MRI slice.
+Jiang et al.[29] proposed CPAConv-POCO, which leverages continuous position-adaptive convolution for lung nodule 3D reconstruction.
+Chen et al.[30] applied completion networks to cartilage repair, using transformer-enhanced point cloud completion to recover femoral cartilage surfaces.
+For anatomical organ-level segmentation, Zhang et al.[31, 32] designed point-voxel fusion networks to segment Couinaud liver segments, showing superior performance on LiTS and 3Dircadb.
+More recently, Yassin et al.[33] proposed a multimodal foundation model for medical point cloud completion based on the large-scale MedShapeNet dataset[13], achieving generalizable completion across 240+ anatomical classes.
+Zhang et al.[34] introduced a hierarchical feature learning framework based on state-space models to further improve anatomical understanding tasks such as completion and segmentation.
+Although point cloud completion shares similarities with upsampling by extracting global features from partial input to complement geometric details, most existing methods are validated on synthetic or restricted datasets.
+In contrast, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling.
+Furthermore, by training on realistic datasets such as MedShapePelvic and evaluating on the Pelvic1k subset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
+II-B Segment Anything
 The Segment Anything Model (SAM)[9] was trained on a dataset of 11 million images and over 1 billion masks.
@@
 By combining its segmentation output with point cloud upsampling, we demonstrate that deep segmentation models can serve as a viable input source for high-fidelity shape reconstruction.
-3 Methodology
-3.1 Overview of the Proposed Framework
-We propose a novel framework called Point Cloud Upsampling as Implicit Statistical Shape Model (PUSSM), which integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction.
-As shown in Figure 1(a), the core idea is to learn an implicit shape prior from real anatomical data through a deep upsampling network, allowing sparse point clouds to be restored to dense representations in an anatomically plausible way.
+III Methodology
+III-A Overview of the Proposed Framework
+We propose a novel framework that integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction.
+As shown in Figure 1(a), the core idea is to learn an implicit anatomical prior from real data through a deep upsampling network, allowing sparse point clouds to be restored to dense and anatomically plausible representations.
 The pipeline consists of two sequential modules.
@@
 Then, as illustrated in Figure 1(b), the sparse point cloud is passed into a pretrained point cloud upsampling network, which has been trained on MedShapePelvic samples.
-Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes.
-Figure 2: Distributional interpretation of implicit statistical shape modeling.
+Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes, yielding smooth triangular meshes suitable for quantitative evaluation and visualization.
+Figure 2: Distributional interpretation of implicit anatomical prioring.
 The segmentation-derived point cloud (left) deviates from the real anatomical distribution.
 Our network $f_{\theta}$ learns to map the sparse inputs toward the ground-truth distribution through point cloud upsampling.
-3.2 Implicit Statistical Shape Modeling
-Unlike classical Statistical Shape Models (SSMs), which explicitly encode anatomical variation via PCA over aligned meshes, our framework learns an implicit statistical shape model through data-driven training.
-Specifically, let $X\subset\mathbb{R}^{3}$ denote a sparse point cloud from SAM-Med3D, and let $Y\subset\mathbb{R}^{3}$ be the corresponding ground truth dense point cloud from MedShapePelvic.
-We train a deep function $f_{\theta}:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{M\times 3}$ to minimize the geometric discrepancy between $f_{\theta}(X)$ and $Y$, using a distance metric $D(\cdot,\cdot)$ such as the Chamfer Distance:
-$\min_{\theta}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[D(f_{\theta}(X),Y)\right]$
-This process embeds the shape prior directly into the network weights $\theta$, enabling the model to generalize to new, noisy, or incomplete segmentation results.
-Thus, $f_{\theta}$ serves as an implicit statistical shape model learned directly from anatomical data—requiring no landmark alignment, shape vectors, or PCA decomposition.
+This reflects an implicit learning of intra-population variation, capturing not only shape completion but also a shift toward the anatomical manifold.
+III-B Learning Implicit Anatomical Priors
+Distributional notation
+Let $X=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{3}$ be the sparse points sampled from SAM-Med3D masks and $Y=\{y_{j}\}_{j=1}^{M}$ the dense target points from MedShapePelvic.
+We regard point clouds as empirical measures $\mu_{X}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ and $\nu_{Y}=\frac{1}{M}\sum_{j=1}^{M}\delta_{y_{j}}$.
+The upsampler $f_{\theta}:\mathbb{R}^{N\times 3}\!\to\!\mathbb{R}^{M\times 3}$ induces a pushforward measure $f_{\theta:}\mu_{X}=\frac{1}{M}\sum_{j=1}^{M}\delta_{\hat{y}_{j}}$ with $\hat{Y}=f_{\theta}(X)=\{\hat{y}_{j}\}_{j=1}^{M}$.
+Learning then amounts to aligning $f_{\theta:}\mu_{X}$ with $\nu_{Y}$ on the anatomical shape manifold approximated by the training set.
+Empirical risk
+We minimize a symmetric Chamfer discrepancy between the predicted and target point sets: Eq. (1)
+$\min_{\theta}\;\mathbb{E}_{(X,Y)\sim\mathcal{D}}\Big[\underbrace{\int\!\!\min_{y\in Y}\|z-y\|_{2}^{2}\,\mathrm{d}\big(f_{\theta:}\mu_{X}\big)(z)}_{\text{pred}\to\text{target}}\;+\;\underbrace{\int\!\!\min_{\hat{y}\in\hat{Y}}\|z-\hat{y}\|_{2}^{2}\,\mathrm{d}\nu_{Y}(z)}_{\text{target}\to\text{pred}}\Big],$ (1)
+which is exactly the usual (squared) Chamfer distance written in a distributional form. In practice, Eq. (1) reduces to
+$\mathrm{CD}(\hat{Y},Y)=\frac{1}{M}\sum_{\hat{y}\in\hat{Y}}\min_{y\in Y}\|\hat{y}-y\|_{2}^{2}+\frac{1}{M}\sum_{y\in Y}\min_{\hat{y}\in\hat{Y}}\|y-\hat{y}\|_{2}^{2}.$
+Optimizing Eq.
+(1) embeds shape regularity and population variability into the network parameters $\theta$, enabling robust densification of sparse, segmentation-derived inputs without landmarks, explicit correspondences, or PCA.
 Figure 2 further illustrates this concept from a distributional perspective, showing how our model maps the segmentation-derived distribution toward the ground-truth anatomical distribution via upsampling.
-Figure 3: Conceptual comparison between classical SSM and our implicit PUSSM framework across key dimensions.
-3.3 Comparison with Classical SSMs
-To highlight the differences between our method and classical SSMs, Figure 3 provides a conceptual comparison across three key dimensions: shape representation, prior data source, and learning strategy.
-Traditional SSMs rely on manually constructed shape vectors and PCA decomposition of landmark-aligned meshes.
-In contrast, our method encodes shape priors directly through end-to-end learning from dense mesh data, enabling greater flexibility and broader applicability.
-3.4 Evaluation Metrics
-3.4.1 For Point Cloud Upsampling
+We observe a consistent reduction in Chamfer Distance variance across test samples after upsampling, indicating convergence to a learned anatomical distribution.
+Ergo, the network aims to reduce geometric variance and improving reconstruction fidelity across test cases.
+III-C Evaluation Metrics
+III-C1 For Point Cloud Upsampling
 To evaluate the point cloud upsampling network, we used three key metrics:
@@
 Given a set of upsampled points $P$ and a reference surface $S$, the Point-to-Surface Distance is defined as:
-[…]
-3.4.2 For Surface Reconstruction
+$\text{P2F}(P,S)=\frac{1}{|P|}\sum_{p\in P}\min_{s\in S}\|p-s\|$
+III-C2 For Surface Reconstruction
 To evaluate the surface reconstruction of our framework, we used three key metrics: Area-Length Ratio, Manifoldness Rate and Connected Component Discrepancy[36, 37].
-Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh.
-This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality.
-It is defined as:
+Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh. This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality.
 Manifoldness Rate (MR) quantifies the proportion of edges in a mesh that satisfy manifold conditions.
@@
 Conversely, a lower rate highlights the presence of non-manifold edges, which may necessitate mesh repair or refinement.
-It is defined as:
-Manifoldness Rate=Number of manifold edgesTotal number of $\text{Manifoldness Rate}=\frac{\text{Number of manifold edges}}{\text{Total number of edges}}$
 Connected Component Discrepancy (CC Diff.) assesses the topological connectedness of a mesh by comparing the observed number of connected components to the expected number.
 A discrepancy of 0 signifies that the mesh has the anticipated number of connected components, implying structural integrity.
-It is defined as:
-[…]
 Those three indicators to evaluate the geometric quality of reconstructed meshes.
 Besides, to evaluate the fidelity of the reconstructed meshes against the ground truth, we employ several widely recognized metrics[38]: Chamfer Distance (CD), F-Score (F1), Normal Consistency (NC), Edge Chamfer Distance (ECD), and Edge F-Score (EF1).
-4 Experiments
-4.1 Settings
-4.1.1 Datasets
-In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[16] as MedShapePlevic for point cloud upsampling to learn shape prior knowledge.
-We then visualized Quantitative analyzed result based on Pelvic1k[19], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
-Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g.
-subdatasets 3, 4, 6).
-For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[16], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[31].
-The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples.
-Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
-4.1.2 Comparison Methods
-In order to verify the effectiveness of our proposed theory, point cloud upsampling is the first part.
-We built a baseline based on MPU[31], PU-GCN[24], Grad-Pu[13] and PUCRN[9].
-In more experiments, only by comparing it with two advanced algorithms: Grad-Pu that realized any ratio upsampling and PUCRN that is the SOTA model.
-For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment.
-And we test the result based on training in MedShapePelvic[16] and PU1k dataset[24], which is a synthetic dataset from 3D models.
-This comparison with identical parameter settings would be presented in ablation: Section 4.3.
-4.1.3 Implementation Details
-Experiments were developed using the PyTorch framework on the Ubuntu 22.04 system.
-We used an NVIDIA L20 GPU with 48GB of graphics memory and an Intel(R) Xeon(R) Platinum 8457C host with 100GB of RAM.
-The network was trained over 100 epochs with a batch size of 64.
-The initial learning rate was set to 0.001, with a decay rate of 0.05.
-4.2 Upsampling Results
-In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShapePelvic dataset in 4×4$\times$4 × upsampling.
-The quantitative results for the dataset in Table 1 show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
-This result demonstrates the effectiveness of PUCRN in CD, HD, and P2F, showcasing its superior capability in learning the shape prior knowledge in pelvic.
-The 3 stage network refinement structure allow it could minimizing nearest-neighbor distance differences and reducing maximum point-to-point distance errors.
-Additionally, noticed that Grad-PU with very high HD.
-This may be caused by the input pelvic training data.
-This reflects that Grad-PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 4 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
-Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2.
-Grad-PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set.
-However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect.
-The results show that PUCRN still has good performance.
-Table 1: Performance comparison of two methods with MedShapePelvic across input scales.
+TABLE I: Performance comparison of two methods with MedShapePelvic across input scales.
 Method 512 input 1024 input 2048 input 4096 input
@@
 PUCRN 3.104 7.806 6.150 3.946 16.635 6.436 1.839 5.709 4.066 1.274 5.087 2.719
-4.3 Ablation Studies
-To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShapePelvic and PU1k dataset[24] have been tested.
-Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples.
-This avoids the difference in learning effectiveness caused by the large difference in the number of training samples.
-The result shows in Table 3
-Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024.
+IV Experiments
+IV-A Settings
+IV-A1 Datasets
+In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[13] as MedShapePelvic for point cloud upsampling to learn shape prior knowledge.
+We then visualized Quantitative analyzed result based on Pelvic1k[39], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
+Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g.
+subdatasets 3, 4, 6).
+For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[13], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[40, 21, 41, 42].
+The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples.
+Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
+IV-A2 Comparison Methods
+In order to verify the effectiveness of our proposed theory, point cloud upsampling is the first part.
+We built a baseline based on MPU[40], PU-GCN[21], Grad-Pu[22] and PUCRN[23].
+In more experiments, only by comparing it with two advanced algorithms: Grad-Pu that realized any ratio upsampling and PUCRN that is the SOTA model.
+For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment.
+And we test the result based on training in MedShapePelvic[13] and PU1k dataset[21], which is a synthetic dataset from 3D models.
+This comparison with identical parameter settings would be presented in ablation: Section IV-C.
+IV-A3 Implementation Details
+Experiments were developed using the PyTorch framework on the Ubuntu 22.04 system.
+We used an NVIDIA L20 GPU with 48GB of graphics memory and an Intel(R) Xeon(R) Platinum 8457C host with 100GB of RAM.
+The network was trained over 100 epochs with a batch size of 64.
+The initial learning rate was set to 0.001, with a decay rate of 0.05.
+TABLE II: Quantitative comparison in $16\times$ upsampling after training in $4\times$. The input points are 512 and 1024.
 Method 512 to 8192 1024 to 16384
@@
 PUCRN 3.586 9.984 6.824 2.870 18.899 5.872
-As a result, training on the MedShapePelvic dataset performed better than training on PU1k.
-This indicates that PUCRN has indeed learned Pelvic’s shape features.
-This is very important.
-However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set.
-These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
-Table 3: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
+TABLE III: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
 Method 512 input 1024 input 2048 input 4096 input
-[…]
+$downarrow CD\downarrow HD\downarrow$ $downarrow CD\downarrow HD\downarrow$
 Grad-PU(MedShapePelvic) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
@@
 PUCRN(PU1k) 3.749 10.065 2.902 9.955 1.995 6.134 1.379 5.253
+IV-B Upsampling Results
+In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShapePelvic dataset in $4\times$ upsampling.
+The quantitative results for the dataset in Table I show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
+This result demonstrates the effectiveness of PUCRN in CD, HD, and P2F, showcasing its superior capability in learning the shape prior knowledge in pelvic.
+The 3 stage network refinement structure allow it could minimizing nearest-neighbor distance differences and reducing maximum point-to-point distance errors.
+Additionally, noticed that Grad-PU with very high HD.
+This may be caused by the input pelvic training data.
+This reflects that Grad-PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 3 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
+Besides, we also conducted $16\times$ upsampling tests, in Table II.
+Grad-PU can perform upsampling at any scale, so we set its ratio directly to $16\times$ on the test set.
+However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two $4\times$ samplings on PUCRN to achieve a $16\times$ sampling effect.
+The results show that PUCRN still has good performance.
+IV-C Ablation Studies
+To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the $4\times$ sampling result based on training in MedShapePelvic and PU1k dataset[21] have been tested.
+Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples.
+This avoids the difference in learning effectiveness caused by the large difference in the number of training samples.
+The result shows in Table III
 Figure 3: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
-4.4 Reconstruction Visualizations
+As a result, training on the MedShapePelvic dataset performed better than training on PU1k.
+This indicates that PUCRN has indeed learned Pelvic’s shape features, which means that it learned statistical variation embedded in anatomical structures, serving as an implicit anatomical prior without relying on PCA or landmarks.
+However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set.
+These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
+TABLE IV: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
+Subdataset 3 Subdataset 4 Subdataset 6
+[…]
+Ground Truth 0.692 100.0 0.26 0.721 100.0 1.68 0.742 100.0 0.66
+Segmented Result (by SAM-Med3D) 0.676 100.0 1.94 0.699 100.0 4.89 0.735 100.0 1.18
+Med-PU (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
+Med-PU (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
+IV-D Reconstruction Visualizations
 In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM.
+TABLE V: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
+Subdataset 3 Subdataset 4 Subdataset 6
+16,384 Points
+by Poisson Sampling
+[…]
+Segmented Result (by SAM-Med3D) 8.379 0.214 0.691 7.420 0.408 56.272 0.126 0.711 54.559 0.250 13.997 0.214 0.716 10.724 0.357
+Med-PU (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
+Med-PU (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
 The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds.
-The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling.
+The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform $16\times$ upsampling.
 The model surface is reconstructed using the Voxel Grid with Marching Cubes[43], where voxel size is $1.5$.
@@
 And there are no obvious contours of layer after reconstructed.
-4.5 Reconstruction Evaluations
-To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in Section3.4.2 For Surface Reconstruction.
-Table 4: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
-Subdataset 3 Subdataset 4 Subdataset 6
-[…]
-Ground Truth 0.692 100.0 0.26 0.721 100.0 1.68 0.742 100.0 0.66
-Segmented Result (by SAM-Med3D) 0.676 100.0 1.94 0.699 100.0 4.89 0.735 100.0 1.18
-PUSSM (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
-PUSSM (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
-As shows in Table 4.
+IV-E Reconstruction Evaluations
+To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in SectionIII-C2 For Surface Reconstruction.
+As shows in Table IV.
 Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy.
@@
 Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation.
-We used metrics (CD, F1, NC, ECD, EF1) in Section3.4.2 For Surface Reconstruction.
+We used metrics (CD, F1, NC, ECD, EF1) in SectionIII-C2 For Surface Reconstruction.
 To ensure a fair and uniform comparison, we employ Poisson disk sampling to generate 16,384 points from each mesh, including the ground truth, segmented results, and upsampled reconstructions.
 This sampling strategy maintains uniform point distribution and adequately captures both global structures and fine details, facilitating reliable metric computation.
-Table 5: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
-Subdataset 3 Subdataset 4 Subdataset 6
-[…]
-Segmented Result (by SAM-Med3D) 8.379 0.214 0.691 7.420 0.408 56.272 0.126 0.711 54.559 0.250 13.997 0.214 0.716 10.724 0.357
-PUSSM (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
-PUSSM (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
-As shows in Table 5.
+As shows in Table V.
 Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations.
@@
 These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions.
-5 Conclusion
-In this work, we proposed PUSSM, an implicit statistical shape modeling framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data.
+And we observe consistent reductions in Chamfer Distance and improvements in Normal Consistency and F-scores, indicating that the upsampling network effectively maps sparse point clouds toward the anatomical shape manifold as Figure 2.
+V Discussion
+This study demonstrates that coupling a robust volumetric segmenter with a dedicated point cloud upsampler can bridge the gap between voxel labels and mesh-quality anatomical models.
+Compared with classical SSMs that rely on landmarks and explicit statistical spaces, our approach learns an implicit prior from realistic pelvic shapes (MedShapePelvic), enabling automatic densification and regularization of sparse, segmentation-derived points into smooth, watertight surfaces suited for visualization and downstream analysis.
+Why domain supervision matters
+Our ablations indicate consistent gains when the upsampler is trained on anatomical meshes (MedShapePelvic) rather than synthetic/CAD-style objects (PU1k).
+This aligns with the intuition that the statistics of cortical thickness, local curvature, and trabecular recesses in pelvic bones differ markedly from generic object datasets.
+Prior CMPB studies similarly show that medical-domain supervision/priors improve realism and downstream usability of the reconstructed geometry [44, 45].
+In practice, domain-aligned supervision helps suppress perforations and over-smoothing, yielding lower HD and improved edge fidelity after meshing.
+Limitations and Future Work
+The performance of our 3D reconstruction pipeline, Med-PU, is limited by several factors.
+First, it depends on the quality of front-end segmentation and is sensitive to imaging domain shifts, such as challenging boundaries or metal artifacts, which may propagate errors to the point set and surface.
+Second, our meshing relies on iso-surface extraction without explicit topology control, and while we apply standard cleanup and report manifoldness-oriented metrics, pathological non-manifold cases may still occur.
+Third, our evaluation is geometry-centric, and the clinical impact—such as interobserver variability in contouring, planning accuracy, or dose/navigation surrogates—remains to be quantified through future user studies.
+Broader evaluations across diverse anatomies, scanners, and acquisition protocols are needed to assess robustness to domain shifts.
+To address inherited segmentation errors, joint or end-to-end training of segmentation and upsampling could be explored.
+Additionally, runtime and memory demands increase with higher upsampling ratios, suggesting the need for lighter backbones, GPU-friendly operators, or uncertainty-aware pruning to enhance efficiency and reliability.
+Finally, integrating task-level endpoints and extending the implicit prior beyond medical applications (e.g., to robotics or industrial inspection) would further validate Med-PU as a general-purpose, knowledge-based 3D reconstruction system.
+Broader surveys, such as those in clinical 3D reconstruction for endoscopy, emphasize the importance of task-driven validation beyond geometric scores [46].
+VI Conclusion
+In this work, we proposed Med-PU, an implicit anatomical prioring framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data.
 Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation.
 Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
-Future work will extend this framework to 3D SAM model[15] by points based prompting to integrate it as Chain-of-Thought (CoT)[12] for Segmentation.
-{credits}
+Instead of explicitly decomposing shape variations, Med-PU learns to embed these variations in a function space.
+This allows the model to restore anatomically plausible shapes from sparse or noisy segmentations, demonstrating its ability to capture and regularize shape diversity.